Sears Appliance Repair has been designed to service one hundred and sixty appliances a day. Taking into account its current operating constraints, Sear’s expects to service eighty appliances per day. Sears has, however, been able to actually service seventy-two appliances per day.

  1. Compute the utilization and efficiency of Sear’s appliance repair center.
  2. Compute the expected output if the efficiency of Sear’s appliance repair center is seventy percent.
Answer #1

a. Utilization = (Actual Output/Design Capacity)*100

= (72/160)*100

= 45%

Efficiency = (Actual Output/Effective Capacity)*100

= (72/80)*100

= 90%

b. If efficiency = 70%

70/100 = (Actual Output/80)

Actual Output = 56 Units